OpenStudy (anonymous):

how to solve 4(3x-y) = 9x + 5 3(2x+3y) = 13(x+y-5)

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@Mashy

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (anonymous):

no the my options are differenrt

OpenStudy (anonymous):

@ememlove

OpenStudy (anonymous):

what are your options?

OpenStudy (anonymous):

(7,4) (10,6) (-3,2) (3,-2)

OpenStudy (anonymous):

@ganeshie8 @Gabylovesyou @kingdeolt

OpenStudy (anonymous):

oh its like that :D I did not realise this was simultaneous equations :D i thought you have to solve to find two equations of a line :P.

OpenStudy (anonymous):

First line simplifies to 3x - 4y - 5 = 0 Second line simplifies to 7x +4y -65 = 0 If you add these together, 10x - 70 = 0 10x = 70 x = 7. Well if x=7 then its your first choice answer :)

OpenStudy (anonymous):

But you can check for y, sub into equation 1 3(7) - 4y - 5 = 0 21-5 = 4y 16 = 4y y = 4 so (7,4)

OpenStudy (anonymous):

im getting -7x

OpenStudy (anonymous):

what do you do and how do you get it?

OpenStudy (anonymous):

I wrote pretty much the whole working out, so check in it,maybe you made a mistake along the way.

OpenStudy (anonymous):

it will be great if u show me the working from the starat

OpenStudy (anonymous):

You mean simplification of equations well in second one 3(2x+3y) = 13(x+y-5) multiply out the brackets 6x + 9y = 13x + 13y - 65 take away 6x and 9y from both sides 7x + 4y - 65 = 0.

OpenStudy (anonymous):

thx was going wrong there